OpenStudy (anonymous):

Solve for x. x/2-x+5/11=6 a.) 142/9 b.) 44/3 c.) 32/9 d.) 142/11

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

assuming it is \[\frac{x}{2}-\frac{x+5}{11}=6\]

OpenStudy (anonymous):

x = 149/2

OpenStudy (anonymous):

x = 142/9